# Tuning the band topology and topological Hall effect in skyrmion crystals via the spin-orbit coupling

### v1

@@ VI Summary
In summary, we presented a comprehensive analysis of the topological Hall effect in the skyrmion crystal in the presence of spin-orbit coupling, of Rashba, Dresselhaus, or the Weyl type. We considered the square lattice as a prototype and the magnetic unit cell contained a skyrmion or an antiskyrmion of either Bloch or Néel type. Detail results for the band structure and the THC were presented for the Rashba SOC, and we showed that the results for the other two cases (Dresselhaus SOC and Weyl SOC) can be expressed in terms of the Rashba results.
The presence of the SOC term affects the electron motion in two ways. First, it modifies the emergent magnetic field seen by the electron as it moves from site to site, and second, the hopping amplitudes are also modified (see Supplementary Materials for a concrete example). While both affect the band dispersion and the THC, the first term modifies the band topology (Chern numbers) as well.
Expressions for the emergent magnetic field were obtained for the isolated skyrmion or antiskyrmion in the presence of all three SOC terms. These results are useful in the interpretation of the electronic band structure and the THC in the skyrmion crystal. The competition between the emerging magnetic field due to the skyrmion texture and the SOC term leads to a tuning of the band dispersion and, therefore, the magnitude of the THC in any type of skyrmion. For example, for the Rashba SOC case and the Néel skyrmion crystal, we showed that for a critical value $\alpha_{R}^{0}$ of the SOC strength, the combined magnetic field due to the skyrmion and the SOC term is uniform enough to produce the band structure similar to that in a uniform magnetic field, leading to prominent Hall plateaus.
@@
We showed that the results for the Dresselhaus SOC and Weyl SOC can be mapped into the case for Rashba SOC, because the Hamiltonians are either the same or they are related by symmetry. Keeping this in mind, we have discussed the Rashba SOC case in detail. Results for the Dresselhaus SOC and Weyl SOC cases, including the THC, can be inferred from the appropriate mapping, which are summarized in Table 1.
Experimenters are increasingly successful in fabricating skyrmions crystals, even though their stability and manipulation hinge on a delicate interplay between magnetic interactions, external fields, and temperature. It is also well established that the strength of the Rashba SOC can be tuned via an external electric field, and also in some cases, experimenters have been able to generate and tune the Dresselhaus SOC via interface engineering. These developments on the experimental front allow for the study of the many interesting phenomena predicted here for the skyrmion crystal including the tuning of the band topology and the THC, as well as the sign reversal of the Hall current and the edge current in the nanoribbons. This could lead to a fertile playground for skyrmion-based spintronics applications.

### v2

@@ VI Summary
In summary, we presented a comprehensive analysis of the topological Hall effect in the skyrmion crystal in the presence of spin-orbit coupling, of Rashba, Dresselhaus, or Weyl type. We considered the square lattice as a prototype and the magnetic unit cell contained a skyrmion or an antiskyrmion of either Bloch or Néel type. Detail results for the band structure and the THC were presented for the Rashba SOC, in terms of which the other two cases can be described.
The presence of the SOC term affects the electron motion in two ways. First, it modifies the emergent magnetic field seen by the electron as it moves from site to site, and second, the hopping amplitudes are also modified. While both affect the band dispersion and the THC, the first term modifies the band topology (Chern numbers) as well.
Expressions for the emergent magnetic field were obtained for the isolated skyrmion or antiskyrmion in the presence of all three SOC terms. These results are useful in the interpretation of the electronic band structure and the THC in the skyrmion crystal. The competition between the emerging magnetic field due to the skyrmion texture and the SOC term leads to a tuning of the band dispersion and, therefore, the magnitude of the THC in any type of skyrmion. For example, for the Rashba SOC case and the Néel skyrmion crystal, we showed that for an optimal value $\alpha_{R}^{0}$ of the SOC strength, the combined magnetic field due to the skyrmion and the SOC term is uniform enough to produce the band structure similar to that in a uniform magnetic field, leading to prominent Hall plateaus.
@@
When the SOC is Dresselhaus or Weyl type, all results can be mapped into the case for Rashba SOC, because the corresponding Hamiltonians are either the same or they are related by symmetry. Keeping this in mind, we have discussed the Rashba SOC case in detail. Results for the Dresselhaus SOC and Weyl SOC cases, including the THC, can be inferred from the appropriate mapping, which are summarized in Table 1.
Experimenters are increasingly successful in fabricating skyrmions crystals, even though their stability and manipulation hinge on a delicate interplay between magnetic interactions, external fields, and temperature. It is also well established that the strength of the Rashba SOC can be tuned via an external electric field, and also in some cases, experimenters have been able to generate and tune the Dresselhaus SOC via interface engineering. These developments on the experimental front allow for the study of the many interesting phenomena predicted here including the tuning of the band topology and the THC, as well as the sign reversal of the Hall current and the edge current in the nanoribbons. This could lead to a fertile playground for skyrmion-based spintronics applications.